Compute the exponentiation x^e mod 29 of x = 5 with both variants of e from above* for n = 4. Use the square-and-multiply algorithm and show each step of your computation.

*above, referring to the formulas e = 2^n + 1 and e = 2^n -1
Answer #1

Algorithm
result = 1;
while(e>0){
   if (e is odd)
result = (result*x) % 29;
e = e/2
x = (x*x) % 29;   
}
print(result)

x^e mod 29 where x= 5 n=4;
method 1
e = 2^n+1 = (2^4+1) = 17;
(1)Here e is odd hence result = (1*5)%29 = 5
e = 17/2 = 8
x = (5*5)%29 = 25

(2)Here e is even
e = 8/2 = 4
x = (25*25)%29 = 625%29 = 16

(3)Here e is even
e = 4/2 = 2
x = (16*16)%29 = 256%29 = 24

(4)Here e is even
e = 2/2 = 1
x = (24*24)%29 = 576%29 = 25

(5)Here e is odd hence result = (5*25)%29 = 125%29 = 9
e = 1/2 =0
Result = 9;

method 2
e = 2^n-1 = (2^4-1) = 15;
(1)Here e is odd hence result = (1*5)%29 = 5
e = 15/2 = 7
x = (5*5)%29 = 25

(2)Here e is odd hence result = (5*25)%29 = 125%29 = 9
e = 7/2 = 3
x = (25*25)%29 = 625%29 = 16

(3)Here e is odd hence result = (9*16)%29 = 144%29 = 28
e = 3/2 = 1
x = (16*16)%29 = 256%29 = 24

(4)Here e is odd hence result = (28*24)%29 = 672%29 = 5
e = 1/2 = 0
Result = 5;